# Isoperimetric inequalities and sharp upper bounds for Aharonov-Bohm eigenvalues on surfaces

### 13 Apr 2026 v1

@@ 1. Introduction
When $\partial M\neq\emptyset$, we consider magnetic Neumann boundary conditions $d^{A}u(N)=0$, that in terms of the dual vector field $\nabla^{A}u:=\nabla u-iuA^{\sharp}$, reads $\langle\nabla^{A}u,N\rangle=0$, where $N$ is the unit outer normal to $\partial M$. Here $A^{\sharp}$ denotes the vector field dual to the $1$-form AA; the expression $\nabla^{A}u$ denotes the magnetic gradient of $u$. Then $\Delta_{A}$ defines a compact self-adjoint operator on $L^{2}(M,dv_{g})$, hence it has a non-negative spectrum made of eigenvalues of finite multiplicity diverging to $\infty$:
@@
where $H^{1}_{A,0}(M)$ is the subspace of $H^{1}_{A}(M)$ of function with zero boundary trace. Also in this case we set $\lambda_{k}(M,g,A)\doteq\lambda_{k}(M,A)$. The magnetic Dirichlet eigenvalues will appear in the proof of the main results.
@@
1.2. Aharonov-Bohm potentials
Often, also singular potential one-forms are allowed, if the singularity is well-behaved. This is the case of Aharonov-Bohm-type potentials. To be more precise, let $\Omega\subset\mathbb{R}^{2}$ be a bounded domain, and $0\in\Omega$. Let
@@ 1.2. Aharonov-Bohm potentials
The $1$-form $A^{(\nu)}_{0}$ is defined in $\mathbb{R}^{2}\setminus\{0\}$, where it is closed but not exact, and has flux $\nu$ around any loop enclosing zero. We consider magnetic Neumann boundary conditions, so that the spectrum is discrete (see e.g., [8]). By what we said above $\mu_{1}(\Omega,A^{(\nu)}_{0})>0$ if and only if $\nu\notin\mathbb{Z}$.
In this paper we are interested in shape optimization results for the first eigenvalue of compact, simply connected surfaces (with or without boundary).
@@ 1.3. State of the art
(1.2) $\mu_{1}(\Omega,A^{(\nu)}_{0})\leq\mu_{1}(D,A^{(\nu)}_{0})$
where $D$ is the disk centered at 0 with the same area: $\left|\Omega\right|=\left|D\right|$. Equality holds if and only if $\Omega=D$. Note that the isoperimetric inequality (1.2) holds for any domain (not necessarily simply connected). The proof uses the classical argument of Weinberger for the second eigenvalue of the Neumann Laplacian [19]. Note that in polar coordinates $(r,\theta)$ around 0, $A^{(\nu)}_{0}=\nu d\theta$.
Spherical domains. On the two-dimensional round sphere $\mathbb{S}^{2}$, fix a point $x_{0}$ and consider polar coordinates $(r,\theta)$ based at $x_{0}$, where $r$ is the geodesic distance from $x_{0}$. Then $A_{x_{0}}^{(\nu)}\doteq\nu d\theta$ defines a closed $1$-form on $\mathbb{S}^{2}$ with flux $\nu$ around any loop enclosing $x_{0}$. Note that $A$ is singular at two points: $x_{0}$ and $x_{0}$. For topogical reasons, we need two punctures to define an Aharonov-Bohm potential on a closed, simply connected surface. We remark that the flux around any loop separating $x_{0}$, $x_{0}$ is equal to $\nu$. The following inequality of isoperimetric type is proved in [8].
@@
Fix $x_{0}\in\mathbb{S}^{2}$ and let $\Omega$ be a smooth, bounded domain containing $x_{0}$ but not $x_{0}$. Let $D$ be the spherical cap centered at $x_{0}$ with the same area: $\left|\Omega\right|=\left|D\right|$. Then, the isoperimetric inequality
@@
In both cases equality holds if and only if $\Omega$ is the spherical cap $D$ centered at $x_{0}$.
Under any of the assumptions a) and b), the isoperimetric inequality is known to hold for the lowest positive eigenvalue of the usual Neumann Laplacian, that is:
@@
We point out a recent growing interest in the eigenvalues of the Neumann Laplacian on spherical domains. In particular, we mention [12] where, in the inequality (1.4) for simply connected domains, the area bound is relaxed; precisely from $|\Omega|\leq 0.5|\mathbb{S}^{2}|$ to $|\Omega|\leq 0.94|\mathbb{S}^{2}|$: in other words, (1.4) holds for simply connected domains with area at most $94\%$ that of the sphere. In the recent paper [4] the authors show analytically that if the hypothesis on simply-connectedness is dropped, then spherical caps are not necessarily maximisers of the second Laplace eigenvalue among domains of fixed area. We mention that the method proposed in this paper for the study of isoperimetric inequalities and eigenvalue upper bounds is different from those of Szegö and Weinberger. In fact, among other results, we prove here that Theorem 1.2, b) holds with no restrictions on $|\Omega|$. Adapting the method to the case of the second Neumann eigenvalue of the standard Laplacian one can prove (1.4) for all simply connected spherical domains with no restrictions on the area. This is done in [16].
@@ 1.4. Aims of the paper
Let us define what is an Aharonov-Bohm potential on a compact, simply connected Riemannian surface $\Omega$ with non-empty boundary $\partial\Omega$. It is determined (up to Gauge equivalence) by the flux $\nu\in\mathbb{R}$ and a pole $z_{0}\in\Omega$. We set, on the punctured domain $\Omega\setminus\{z_{0}\}$:
@@
and $\star$ is the Hodge-star operator associated with the Riemannian metric. Note that $\nu$ is the flux of $A^{(\nu)}_{z_{0}}$ around $\partial\Omega$ or any loop enclosing $z_{0}$.
If we have a closed, simply connected Riemannian surface $\Sigma$, an Aharonov-Bohm potential is determined by the flux $\nu$ and two poles $z_{1},z_{2}$. We set, on $\Sigma\setminus\{z_{1},z_{2}\}$:
@@
In this case, $\nu$ is the flux of $A^{(\nu)}_{z_{1},z_{2}}$ around any loop separating the poles. Here $\delta_{P}$ denotes the Dirac delta at a point $P$.
As we already mentioned, in the case of a geodesic disk in $\mathbb{S}^{2}$ with pole $x_{0}$ in its center, $A^{(\nu)}_{x_{0}}$ takes the simpler form $A^{(\nu)}_{z_{0}}=\nu d\theta$, where $(r,\theta)$ are the standard polar coordinates centered at $x_{0}$ (with $r$ the distance from $x_{0}$). This is also the form of the Aharonov-Bohm potential on the standard sphere $\mathbb{S}^{2}$ with poles $x_{0}$, $x_{0}$.
@@
The inequalities are obtained by employing the method of prescribed level lines which could be dated back to Rayleigh and Ritz (see e.g., [14]), and applying it to the Green function of $\Omega$ or $\Sigma$ as above. The same method, using test functions which are constant on the level lines of the torsion function, has been used in [7] to establish a reverse Faber-Krahn inequality for the first eigenvalue of the magnetic Laplacian with constant magnetic field on simply connected planar domains.
First, we consider the isoperimetric problem and prove that among all simply connected surfaces with boundary, gaussian curvature bounded above by $1$, and fixed area smaller than $|\mathbb{S}^{2}|=4\pi$, the first eigenvalue is maximal for a geodesic disk in $\mathbb{S}^{2}$ with the pole of the potential at its center. In particular, among all simply connected spherical domains of fixed area, the geodesic disk with pole at the center is maximal. Note that we have no restriction on the area, extending in this way the result of [8], in particular Theorem 1.2, b).
Another natural question is the following. Consider all simply connected, closed Riemannian surfaces of fixed area: does there exist a maximiser for the first eigenvalue? For the second eigenvalue of the usual Laplacian the answer is positive: it is the round sphere $\mathbb{S}^{2}$ (if we fix the area to be $4\pi$), and the result is due to Hersch [10]. Quite surprisingly, in the magnetic case there is no maximiser for $\mu_{1}(\Sigma,A^{(\nu)}_{z_{1},z_{2}})$ for fixed area, see Subsection 4.7. In this paper we prove an upper bound which is sharp and depends only on an upper bound on the gaussian curvature, see Theorem 2.3.
If the surface is $\mathbb{S}^{2}$ with two poles, a natural question is then: what is the optimal position of the two poles, for which fhe first eigenvalue is maximal? We show that the two poles must be antipodal.
Even if the focus of the paper is on simply connected surfaces, with or without boundary, we study here also the case of Riemannian annuli (surfaces diffeomorphic to $\mathbb{S}^{1}\times[0,1]$). Then we are able to prove that the first magnetic eigenvalue is maximal for the unique flat cylinder in a given conformal class.
@@ 1.5. Organization of the paper
The paper is organized as follows: in Section 2 we state the main results, namely Theorem 2.1 (isoperimetric inequality) and 2.3 (sharp upper bound) and their consequences on spherical domains, namely Corollaries 2.2 and 2.4. In Section 3 we prove Theorem 2.1. In Section 4 we prove Theorem 2.3 and discuss its sharpness. We have included two appendices. In Appendix A we compute explicitly the magnetic spectrum of the sphere with opposite poles and discuss other properties of eigenvalues on spherical domains. In Appendix B we discuss the case of Riemannian annuli.
@@ 2.1. Isoperimetric inequality for simply connected surfaces with boundary
For a constant $K>0$, we let $\mathcal{M}_{K}$ denote the family of compact simply connected Riemannian surfaces with smooth boundary and gaussian curvature bounded above by $K$. The model domains for our eigenvalue problem will be spherical caps in $\mathbb{S}^{2}_{K}$, the round sphere with constant curvature $K$.
@@ 2.2. Upper bounds for simply connected surfaces
We can also consider compact, simply connected surfaces with boundary and area larger than $|\mathbb{S}^{2}_{K}|$. An analogous upper bound holds.
Theorem 2.5.
Let $\Omega\in{\mathcal{M}}_{K}$, let $z_{0}\in\Sigma$ and let $A_{z_{0}}^{(\nu)}$ be an Aharonov-Bohm potential with flux $\nu\in(0\frac{1}{2}]$ and pole $z_{0}$. Then we have:
(2.2) $\mu_{1}(\Sigma,A_{z_{0}}^{(\nu)})$ $\leq K\nu(\nu+1)$
$=\mu_{1}(\mathbb{S}^{2}_{K},A_{x_{0}}^{(\nu)})$
@@ 2.3. Riemannian annuli
Finally, we briefly consider the case of a Riemannian annulus $\mathcal{C}$, namely, a compact surface diffeomorphic to $\mathbb{S}^{1}\times[0,1]$ endowed with a Riemannian metric. Let $A^{(\nu)}$ be a smooth closed $1$-form on $\Sigma$ with flux $\nu\in\mathbb{R}$ around one (each) of the boundary circles. For example, in cylindrical coordinates $(\theta,z)$, one can take $A^{(\nu)}=\nu d\theta$. We consider the magnetic Laplacian $\Delta_{A^{(\nu)}}$. Then the Neumann spectrum is discrete, and the first eigenvalue is positive as long as $\nu\notin\mathbb{Z}$ and one has by the min-max principle
@@
Now, any Riemannian annulus $\mathcal{C}$ is conformally equivalent to a unique flat cylinder $\mathcal{C}_{M}\doteq\mathbb{S}^{1}\times[-M,M]$, and $M$ is the so-called conformal modulus of $\mathcal{C}$. We refer e.g., to [15] where the above study is carried out for the magnetic Steklov problem. Here is the relevant result:
@@ 3.1. Reduction to a one dimensional problem
Let $\nu\in(0,\frac{1}{2}]$ and let $z_{0}\in\Omega$ be fixed. Let $\psi_{z_{0}}$ be the Green function on $\Omega$ with pole $z_{0}$ defined in (1.5). Recall that
@@
As we shall see, the space $\mathcal{R}(\Omega)$ consists of all functions in $H^{1}_{A}(\Omega)$ that are constant on the level lines of $\psi$. We set
@@
and this follows from the geometric isoperimetric inequality (3.9), valid for spherical domains (see (3.10)).
@@
Let $u$ be real-valued and of the form $u=g\circ\psi$. Then
$|d^{A}u|^{2}=(g^{\prime}\circ\psi)^{2}|d\psi|^{2}+(g\circ\psi)^{2}|A|^{2}=\left((g^{\prime}\circ\psi)^{2}+(g\circ\psi)^{2}\right)|d\psi|^{2}$
@@
$\int_{\Omega}|d^{A}u|^{2}=\int_{0}^{\infty}\left(g^{\prime}(t)^{2}+4\pi^{2}\nu^{2}g(t)^{2}\right)\int_{\psi=t}|d\psi|\,dt=\int_{0}^{\infty}\left(g^{\prime}(t)^{2}+4\pi^{2}\nu^{2}g(t)^{2}\right)dt$
where the last identity follows since
@@
Recall that $\psi$ has no critical points in $\overline{\Omega}$, hence $\alpha:(0,\infty)\to(0,M)$ is smooth, strictly decreasing, and admits a smooth inverse $\beta:(0,M)\to(0,\infty)$. We change variable and set

### 29 Apr 2026 v2

@@ 1. Introduction
When $\partial M\neq\emptyset$, we consider magnetic Neumann boundary conditions $d^{A}u(N)=0$, which in terms of the dual vector field $\nabla^{A}u:=\nabla u-iuA^{\sharp}$, read $\langle\nabla^{A}u,N\rangle=0$, where $N$ is the unit outer normal to $\partial M$. Here $A^{\sharp}$ denotes the vector field dual to the $1$-form AA; the expression $\nabla^{A}u$ denotes the magnetic gradient of $u$. Then $\Delta_{A}$ defines a compact self-adjoint operator on $L^{2}(M,dv_{g})$, hence it has a non-negative spectrum made of eigenvalues of finite multiplicity diverging to $\infty$:
@@
where $H^{1}_{0,A}(M)$ is the subspace of $H^{1}_{A}(M)$ of functions with zero boundary trace. Also in this case we set $\lambda_{k}(M,g,A)\doteq\lambda_{k}(M,A)$. The magnetic Dirichlet eigenvalues will appear in the proof of the main results.
@@ 1.2. Aharonov-Bohm potentials
On a simply connected domain $\Omega$, an Aharonov-Bohm potential with pole at $x_{0}\in\Omega$ is a closed smooth $1$-form on $\Omega\setminus\{x_{0}\}$. The circulation of $A$ around $\partial\Omega$ (and around any loop enclosing $x_{0}$) is a real number $\nu$ which is called the flux of $A$.
In $\mathbb{R}^{2}$, an example of Aharonov-Bohm potential with pole at the origin and flux $\nu$ is
@@
We consider magnetic Neumann boundary conditions, so that the spectrum is discrete (see e.g., [8]). By what we said above $\mu_{1}(\Omega,A^{(\nu)}_{0})>0$ if and only if $\nu\notin\mathbb{Z}$.
We remark that any two Aharonov-Bohm potentials with pole at $x_{0}$ and flux $\nu$ differ by an exact form, hence give rise to the same spectrum.
In this paper we are interested in shape optimization results for the first Aharonov-Bohm eigenvalue of compact, simply connected surfaces (with or without boundary).
@@ 1.3. State of the art
(1.2) $\mu_{1}(\Omega,A^{(\nu)}_{x_{0}})\leq\mu_{1}(D,A^{(\nu)}_{0})$
for all $x_{0}\in\Omega$. Here $D$ is the disk centered at 0 with the same area: $\left|\Omega\right|=\left|D\right|$. Equality holds if and only if $\Omega=D$ and $x_{0}=0$. In fact, the isoperimetric inequality (1.2) holds for any domain (not necessarily simply connected). The proof uses the classical argument of Weinberger for the second eigenvalue of the Neumann Laplacian [19]. Note that in polar coordinates $(r,\theta)$ around 0, $A^{(\nu)}_{0}=\nu d\theta$. The result holds also for bounded, smooth domains of the hyperbolic plane $\mathbb{H}^{2}$, with the same proof.
Spherical domains. On the two-dimensional round sphere $\mathbb{S}^{2}$, fix a point $x_{0}$ and consider polar coordinates $(r,\theta)$ based at $x_{0}$, where $r$ is the geodesic distance from $x_{0}$. Then $A_{x_{0}}^{(\nu)}\doteq\nu d\theta$ defines a smooth $1$-form on $\mathbb{S}^{2}\setminus\{x_{0},-x_{0}\}$ which is singular at the two poles: $x_{0}$ and $x_{0}$, and defines a Aharonov-Bohm potential on $\mathbb{S}^{2}$. We note that the flux around any loop that separates $x_{0}$, $x_{0}$ is equal to $\nu$. The following inequality of isoperimetric type is proved in [8].
@@
Fix $x_{0}\in\mathbb{S}^{2}$ and let $\Omega$ be a smooth bounded domain that contains $x_{0}$ but not $x_{0}$. Let $D$ be the spherical cap centered at $x_{0}$ with the same area: $\left|\Omega\right|=\left|D\right|$. Then, the isoperimetric inequality
@@
If $\nu\notin\mathbb{Z}$, equality holds in a) and b) if and only if $\Omega$ is the spherical cap $D$ centered at $x_{0}$.
We remark that under any of the assumptions a) and b), the isoperimetric inequality is known to hold for the lowest positive eigenvalue of the usual Neumann Laplacian, that is:
@@
We point out a recent growing interest in the eigenvalues of the Neumann Laplacian on spherical domains. In particular, we mention [12] where, in the inequality (1.4) for simply connected domains, the area bound is relaxed; precisely from $|\Omega|\leq 0.5|\mathbb{S}^{2}|$ to $|\Omega|\leq\frac{16}{17}|\mathbb{S}^{2}|$: in other words, (1.4) holds for simply connected domains with area at most $94\%$ that of the sphere. In the recent paper [4] the authors show that if the hypothesis on simply-connectedness is dropped, then spherical caps are not necessarily maximisers of the second Laplace eigenvalue among domains of fixed area.
The approach we propose in this paper for the study of isoperimetric inequalities and eigenvalue upper bounds is different from those of Szegö and Weinberger. In fact, among other results, we prove here that Theorem 1.2, b) holds with no restrictions on $|\Omega|$. We remark that adapting the method to the case of the second Neumann eigenvalue of the standard Laplacian one can prove (1.4) for all simply connected spherical domains, with no restrictions on the area. This is done in [16].
@@ 1.4. Aims of the paper
Let us give an analytical description of an Aharonov-Bohm potential on a compact, simply connected Riemannian surface $\Omega$ with non-empty boundary $\partial\Omega$ and pole $z_{0}$. Consider, on the punctured domain $\Omega\setminus\{z_{0}\}$ the smooth $1$-form
@@
and $\star$ is the Hodge-star operator associated with the Riemannian metric. Note that $A_{z_{0}}^{(\nu)}$ is closed and has flux $\nu$ around $\partial\Omega$ (and around any loop enclosing $z_{0}$).
On the other hand, if $\Sigma$ is a compact, simply connected Riemannian surface with empty boundary, an Aharonov-Bohm potential is determined by the flux $\nu$ and two poles $z_{1},z_{2}$. We set, on $\Sigma\setminus\{z_{1},z_{2}\}$:
@@
In this case, $\nu$ is the flux of $A^{(\nu)}_{z_{1},z_{2}}$ around any loop separating the poles and $\delta_{P}$ denotes the Dirac delta measure at a point $P$.
As we already mentioned, in the case of a geodesic disk in $\mathbb{S}^{2}$ with pole $x_{0}$ in its center, $A^{(\nu)}_{x_{0}}$ takes the simpler form $A^{(\nu)}_{z_{0}}=\nu d\theta$, where $(r,\theta)$ are the standard polar coordinates centered at $x_{0}$ (with $r$ being the distance from $x_{0}$). This is also the form of the Aharonov-Bohm potential on the standard sphere $\mathbb{S}^{2}$ with antipodal poles $x_{0}$, $x_{0}$.
@@
The inequalities are obtained by employing the method of prescribed level lines which could be dated back to Rayleigh and Ritz (see e.g., [14]), and by applying it to the Green function of $\Omega$ or $\Sigma$ as above. The same method, using test functions which are constant on the level lines of the torsion function, has been used in [7] to establish a reverse Faber-Krahn inequality for the first eigenvalue of the magnetic Laplacian with constant magnetic field on simply connected planar domains, in the weak field regime.
First, we consider the isoperimetric problem and prove that among all simply connected surfaces with boundary, having gaussian curvature bounded above by $1$ and fixed area smaller than $|\mathbb{S}^{2}|=4\pi$, the first eigenvalue is maximal for a geodesic disk in $\mathbb{S}^{2}$ with the pole of the potential at its center. In particular, among all simply connected spherical domains of fixed area, the geodesic disk with pole at the center is maximal. Note that we have no restriction on the area, extending in this way the result of [8], in particular Theorem 1.2, b).
Another natural question is the following. Consider all simply connected, closed (i.e. compact with empty boundary) Riemannian surfaces of fixed area punctured at two points: does there exist a maximiser for the first Aharonov-Bohm eigenvalue? For the second eigenvalue of the usual Laplacian the answer is positive: it is the round sphere $\mathbb{S}^{2}$, and the result is due to Hersch [10], and can be conveniently stated in terms of the lowest normalized eigenvalue, which is invariant by homotheties:
$\sup_{\Sigma}\left|\Sigma\right|\mu_{2}(\Sigma)=\left|\mathbb{S}^{2}\right|\mu_{2}(\mathbb{S}^{2})=8\pi.$
A bit surprisingly, in the magnetic case there is no maximiser for $\left|\Sigma\right|\mu_{1}(\Sigma,A^{(\nu)}_{z_{1},z_{2}})$; in fact, if $\nu\notin\mathbb{Z}$:
$\sup_{\begin{subarray}{c}\Sigma\end{subarray},\,{z_{1},z_{2}\in\Sigma}}\left|\Sigma\right|\mu_{1}(\Sigma,A^{(\nu)}_{z_{1},z_{2}})$$\infty$,
see Theorem 4.5. The point is that when the flux is not an integer, the first eigenvalue depends on the positions of the poles. When the flux is an integer, the position of the poles have no influence on the spectrum, by gauge invariance. In this paper we prove an upper bound which depends only on an upper bound on the gaussian curvature $K$, see Theorem 2.4. The dependence on $K$ is sharp.
If the surface is $\mathbb{S}^{2}$ punctured at two poles, a natural question is then: what is the optimal position of the poles, for which the first eigenvalue is maximal? We show in Corollary 2.5 that the two poles must be antipodal (in the appendix we compute explicitly the spectrum in that case).
Even if the focus of the paper is on simply connected surfaces, with or without boundary, we study here also the case of Riemannian annuli (surfaces diffeomorphic to $\mathbb{S}^{1}\times[0,1]$). Then we are able to prove that the first Aharonov-Bohm eigenvalue is maximal for the unique flat cylinder in a given conformal class.
@@ 1.5. Organization of the paper
The paper is organized as follows: in Section 2 we state the main results, namely Theorem 2.1 (isoperimetric inequality) and 2.4 (sharp upper bound) and their consequences on spherical domains, namely Corollaries 2.2 and 2.5. In Section 3 we prove Theorem 2.1. In Section 4 we prove Theorem 2.4 and discuss its sharpness. We have included two appendices. In Appendix A we compute explicitly the magnetic spectrum of the sphere with opposite poles and discuss other properties of eigenvalues on spherical domains. In Appendix B we discuss the case of Riemannian annuli.
@@ 2.1. Isoperimetric inequality for simply connected surfaces with boundary
For a constant $K>0$, we let $\mathcal{M}_{K}$ denote the family of compact simply connected Riemannian surfaces with smooth boundary and gaussian curvature bounded above by $K$. When $\left|\Omega\right|\leq\left|\mathbb{S}^{2}_{K}\right|$ the model domain for our eigenvalue problem will be a spherical cap in $\mathbb{S}^{2}_{K}$ with same area (eventually, the whole sphere when $\left|\Omega\right|=\left|\mathbb{S}^{2}_{K}\right|$). We get in that case:
@@
This inequality was first proved with the restriction $|\Omega|\leq 2\pi$ in [8].
When the domain $\Omega\in\mathcal{M}_{K}$ has area greater than that of $\mathbb{S}^{2}_{K}$ the model domain will be the whole sphere, in the following sense.
Theorem 2.3.
Let $\Omega\in{\mathcal{M}}_{K}$, let $z_{0}\in\Sigma$ and let $A_{z_{0}}^{(\nu)}$ be an Aharonov-Bohm potential with flux $\nu\in(0,\frac{1}{2}]$ and pole $z_{0}$. Assume that $\left|\Omega\right|\geq\left|\mathbb{S}^{2}_{K}\right|$. Then we have:
(2.1) $\mu_{1}(\Sigma,A_{z_{0}}^{(\nu)})$ $\leq K\nu(\nu+1)$
$=\mu_{1}(\mathbb{S}^{2}_{K},A_{x_{0}}^{(\nu)})$
@@ 2.3. Riemannian annuli
Finally, we briefly consider the case of a Riemannian annulus $\mathcal{C}$, namely, a compact surface diffeomorphic to $\mathbb{S}^{1}\times[0,1]$ endowed with a Riemannian metric. Let $A^{(\nu)}$ be a smooth closed $1$-form on $\Sigma$ with flux $\nu\in\mathbb{R}$ around any of the boundary circles. For example, in cylindrical coordinates $(\theta,z)$, one can take $A^{(\nu)}=\nu d\theta$. We consider the magnetic Laplacian $\Delta_{A^{(\nu)}}$. Then the Neumann spectrum is discrete, and the first eigenvalue is positive as long as $\nu\notin\mathbb{Z}$ and one has by the min-max principle
@@
Now, any Riemannian annulus $\mathcal{C}$ is conformally equivalent to the flat cylinder $\mathcal{C}_{M}\doteq\mathbb{S}^{1}\times[-M,M]$ for a unique $M>0$; the number $M$ is the so-called conformal modulus of $\mathcal{C}$. We refer e.g., to [15] where the above study is carried out for the magnetic Steklov problem. Here is the relevant result:
@@ 3.1. Reduction to a one dimensional problem
Let $\nu\in(0,\frac{1}{2}]$ and let $z_{0}\in\Omega$ be fixed. Let $\psi_{z_{0}}$ be the Green function in $\Omega$ with pole $z_{0}$ defined in (1.5). Recall that
@@
The space $\mathcal{R}(\Omega)$ consists of all functions in $H^{1}_{A}(\Omega)$ that are constant on the level lines of $\psi$. We set
@@
which follows from the geometric isoperimetric inequality (3.9), valid for spherical domains (see (3.10)).
@@
Let $u$ be real-valued and of the form $u=g\circ\psi$. Then, since $|A|^{2}=4\pi^{2}\nu^{2}|d\psi|^{2}$:
$|d^{A}u|^{2}=(g^{\prime}\circ\psi)^{2}|d\psi|^{2}+(g\circ\psi)^{2}|A|^{2}=\left((g^{\prime}\circ\psi)^{2}+4\pi^{2}\nu^{2}(g\circ\psi)^{2}\right)|d\psi|^{2}$
@@
$\int_{\Omega}|d^{A}u|^{2}=\int_{0}^{\infty}\left(g^{\prime}(t)^{2}+4\pi^{2}\nu^{2}g(t)^{2}\right)\int_{\psi=t}|d\psi|\,dt=\int_{0}^{\infty}\left(g^{\prime}(t)^{2}+4\pi^{2}\nu^{2}g(t)^{2}\right)dt.$
The last identity follows since
@@
Recall that $\psi$ has no critical points in $\overline{\Omega}\setminus\{z_{0}\}$, hence $\alpha:(0,\infty)\to(0,M)$ is smooth, strictly decreasing, and admits a smooth inverse $\beta:(0,M)\to(0,\infty)$. We change variable and set